Reports 1-1 of 1 Clear search Modify search
MIF (General)
shinji.miyoki - 8:20 Saturday 01 February 2025 (32540) Print this report
Comment to OMC DCPD B bad condition (32535)

Consequently, the countermeasures for protecting DCPD from burning could not function under the PRFPMI+DC with 5W injection as before.

 

Accident reason

The strong laser beam from the resonated OMC could not be quenched quickly because of the LPF in the HV driver for OMC PZT (circuit) even if the input signal to this PZT driver was quickly terminated and swung with triangular shape signals in the DGS control system.

You can find the LPF consisting of R151(50K) and C111(0.47uF) and C of PZT(~0.5uF according to speck) to "hv_omc". The cut frequency is ~ 3.2Hz which is much lower than the frequency with which we would like to cut signals.

We need to check whether there is another possibility to slow the signal response to PZT.

 

One of Countermeasures

Simply, the cut of frequency of LPF should be enlarged. However, it will result in the noise introduction of the HV amp for the high-frequency ranges. Aso-kun will ask Arai-san in LIGO what is their methods. 

 

Plan for several weeks

  • Laser power of 1W at the IMC out is kept.
  • Realize the stable operation of PRFPMI+RF with 10W injection beam because we encounter the yaw motion oscillation maybe due to the radiation pressure in FP cavities.
  • DAC noise reduction around 100Hz.
  • Total test of the DCPD protection with safe laser power
    • Firstly, we measure the OMC output changes when the OMC resonance is cut by using one arm FP IR resonance and OMC lock.
    • Modify the circuit if we have a reasonable plan.
    • We try the same measurement and check the speed of changes of the OMC output power.

 

When we will repair the DCPD

  • It depends on how fast the HV circuit can be modified.
  • The time limit will be the end of February when the OMC black shroud will be prepared.

 

Search Help
×

Warning

×